OpenStudy (sh3lsh):

what is the polar form of (7-7i)

OpenStudy (anonymous):

real x axis and nonreal on y axis

OpenStudy (anonymous):

sorry imaginary on y axis

OpenStudy (anonymous):



OpenStudy (anonymous):

I think that's what they are asking. Correct me if I'm wrong. Not 100% sure

OpenStudy (sh3lsh):

well the answer choices are 7sqrt2(cos(7pi/4+ i sin(7pi/4) and stuff like that

OpenStudy (mani_jha):

Find the magnitude. Let z=7-7i \[\left| z \right|=\sqrt{7^{2}+7^{2}}\] \[=7\sqrt{2}\] Find the argument. \[\tan \theta=-7/7=-1\] \[\theta=2\pi+\pi/4\]=\[7\pi/4\] The answer follows : \[z=\left| z \right|(\cos \theta+i \sin \theta)\]\[=\left| z \right| e ^{i \theta}\]

OpenStudy (anonymous):